OpenStudy (anonymous):

What is the value of m? 14+2m=4m-16

pooja195 (pooja195):

14+2m=4m−16 14+2m=4m+−16 2m+14=4m−16 Subtract 4m from both sides. 2m+14−4m=4m−16−4m −2m+14=−16 Subtract 14 from both sides. −2m+14−14=−16−14 −2m=−30 Divide both sides by -2 to get your answer m=?

pooja195 (pooja195):

@madhu.mukherjee.946 sorry but yr wrong.

OpenStudy (madhu.mukherjee.946):